OpenStudy (anonymous):

2x/5y=6, what is the value of y in terms of x??

OpenStudy (anonymous):

\[2*x/5*y=6\]

OpenStudy (anonymous):

oops, \[(2*x)/(5*y)=6\]

OpenStudy (anonymous):

ok, you can multiply both sides by 5y therefore \[(2*x)/(5*y)=6\] becomes \[(5*y)*(2*x)/(5*y)=6*(5*y)\]

OpenStudy (anonymous):

5y cancels on both sides \[2*x = 6*(5*y)\]

OpenStudy (anonymous):

now isolate y

OpenStudy (anonymous):

by dividing both sides by 30, the result of 6*5

OpenStudy (anonymous):

what about the 2?

OpenStudy (anonymous):

you keep the 2 you then have 2/30 * x = y

OpenStudy (anonymous):

so 1/15x=y??

OpenStudy (anonymous):

Multiply through by 5y 2x = 30y -> y = 2x/30 = x/15